Locate and classify the stationary points of the function

f( x,y ) x^2+y^2–6xy +6x +3y –4
1
Expert's answer
2020-03-10T12:04:56-0400

"f_x'=2x-6y+6" ,"f_y'=2y-6x+3" , this implies "x=\frac{15}{16},y=\frac{21}{16}" is a stationary point.

"f_{xx}''=2","f_{xy}''=-6" ,"f_{xy}''=2" hence "(f_{xy}'')^2-(f_{xx}'')(f_{yy}'')^2>0", hence there is no extremum at this point, it is a saddle point.